### v1

# Generalized Euler Equation from Effective Action: Implications for the Smarr Formula in AdS Black Holes

Abstract
We derived a generalized Euler equation, $\epsilon+p=sT+\mu q+y\frac{\partial p}{\partial y}$, using the effective field theory formulation of perfect fluids. This generalization was achieved by introducing a new variable $y$ into the effective action, which encodes geometrical information about the spacetime where the fluid is on. Notably, the generalized Euler equation is independent of the AdS/CFT correspondence. However, when applied to a holographic perfect fluid, this equation naturally recovers the Smarr formula for AdS black holes, thus situating the physical interpretation of the Smarr formula within the framework of Gauge/Gravity duality. Finally, our findings raise questions regarding the validity of the extended black hole thermodynamics proposal, which associates the cosmological constant $\Lambda$ with a thermodynamic variable.
1 Introduction
The AdS/CFT correspondence Juan1 has greatly advanced our understanding of quantum field theory in the strong coupling regime, in part by geometrizing abstract concepts from quantum physics. For instance, a thermal state in the field theory is dual to an eternal AdS black hole Juan2 . Moreover, the Hawking-Page transition between thermal AdS spacetime and a "large" AdS black hole Hawking is interpreted as the confinement/deconfinement transition in a gauge theory on $S^{1}\times S^{3}$ Witten . Within this framework, Karch derived the Smarr formula for AdS black holes from a holographic perspective.
In this article, we revisit the physical meaning of the Smarr formula within the context of gauge/gravity duality. Our goal is to explore the Smarr formula from the perspective of fluid/gravity duality Hubeny1 ; Ranga ; Hubeny2 . It is well-established that the boundary stress tensor of an AdS black hole takes the form of a perfect fluid in global equilibrium Clifford0 ; Clifford . From this, we observe that the thermodynamics of the holographic fluid is, in essence, the thermodynamics of black holes. This opens a bridge between the Smarr formula and the theory of perfect fluids that will investigate.
Another important development relevant to this work is the recent construction of effective actions for perfect fluids Dubovsky1 ; Son ; Dubovsky2 ; Liu2 ; Liu3 ; Ranga1 ; Ranga2 ; Ranga3 ; Ranga4 ; Jensen1 . In this framework, the effective action simultaneously reproduces the perfect fluid stress tensor and ensures the validity of the Euler equation, as the effective Lagrangian functions as a local thermodynamic potential. As we will discuss in Section 2, when the Smarr formula is expressed in a suitable form, it resembles the Euler equation with an additional correction. We interpret this as an indication of the existence of a generalization of the Euler equation, insight that we confirm by deriving a generalized Euler equation from the effective action for perfect fluids. This is achieved by introducing a new variable into the local thermodynamic potential, which is both diffeomorphism-invariant with respect to spacetime transformations and gauge-invariant under the fluid’s emergent low-energy gauge transformations. As we will demonstrate, this new variable has a geometrical interpretation and leads to the central result of this article:
$\epsilon+p=sT+\mu q+y\frac{\partial p}{\partial y}~{}.$ (1)
We refer to this as the generalized Euler equation, with the additional term $y\frac{\partial p}{\partial y}$ representing an extra contribution mentioned earlier. Importantly, this result is independent of black hole physics and the AdS/CFT correspondence. We then apply this equation to the case of a holographic perfect fluid, from which the Smarr formula is naturally reproduced.
Another context in which the Smarr formula has been extensively discussed is extended black hole thermodynamics (EBHT). In essence, this framework offers a specific interpretation of the Smarr formula by introducing a thermodynamic variable associated with the cosmological constant $\Lambda$ Kastor ; Gibbons ; Dolan1 ; Dolan2 ; Dolan3 . Initially, this new thermodynamic variable was identified as a pressure term $p_{\Lambda}\sim\frac{1}{l^{2}}$, but in more recent versions of EBHT, it has been reinterpreted as the central charge $C\sim N^{2}$ where $N$ represents the number of degree of freedom of the dual field theory Visser1 ; Gao ; Visser2 ; Mann6 . This revision was motivated by the work in Karch , which sought to make EBHT consistent with holography. However, the critical point for this article is that our derivation of the generalized Euler equation—and its physical interpretation—conflicts with the interpretation of the Smarr formula as proposed by EBHT. More detailed comments on this issue will be provided in Section 5.
This paper is organized as follows: In Section 2, we review the Smarr formula and highlight its preliminary similarities with the Euler equation. In Section 3, we examine the standard construction of the effective action for perfect fluids and introduce a new variable into the effective action, which enables us to derive a generalized Euler equation. In Section 4, we demonstrate that the Smarr formula is a specific instance of the generalized Euler equation. Section 5 discusses why our results contradict the proposal for extended black hole thermodynamics. Finally, in Section 6, we present our conclusions and suggest potential directions for future research.
2 Smarr formula for AdS black holes
In this section, we will review the Smarr formula for the AdSd Reissner-Nordström black hole, closely following the conventions in Karch . The Einstein-Maxwell action, with a negative cosmological constant, is given by
$S=\frac{1}{16\pi G_{N}}\int d^{d}x\sqrt{-g}\left(R-2\Lambda-F^{2}\right)~{},$ (2)
where $G_{N}$ is the $d$-dimensional Newton’s constant. The negative cosmological constant $\Lambda$ can be expressed in terms of the AdS radius $l$ as follows: $\Lambda=-\frac{(d-2)(d-1)}{2l^{2}}$. For the spherical AdSd Reissner-Nordström black hole, the metric and gauge field are, respectively,
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}d\Omega_{d-2}^{2}~{},\quad f(r)=\frac{r^{2}}{l^{2}}+1-\frac{\tilde{m}}{r^{d-3}}+\frac{\tilde{q}^{2}}{r^{2d-6}}~{},$ (3)
$A=\left(\mu l-\frac{1}{\gamma}\frac{\tilde{q}}{r^{d-3}}\right)dt~{},\quad\gamma=\sqrt{\frac{2(d-3)}{(d-2)}}~{}.$ (4)
The constant $\mu$ is eventually identified with the chemical potential, and it is defined such that the gauge field $A$ vanishes at the event horizon located at $r_{h}$
$\mu=\frac{1}{\gamma}\frac{\tilde{q}}{lr^{d-3}_{h}}~{}.$ (5)
The black hole energy $M$ and charge $Q$ are related to the parameters $\tilde{m}$ and $\tilde{q}$ as follows:
$M=\frac{(d-2)A_{d-2}\tilde{m}}{16\pi G_{N}}~{},\quad Q=\frac{l\gamma(d-2)A_{d-2}\tilde{q}}{8\pi G_{N}}~{},$ (6)
where $A_{D}$ is the area of the $D$-dimensional unit sphere $S^{D}$. The black hole entropy $S$ and the Hawking temperature $T$ are given, respectively, by
$S=\frac{A_{d-2}r_{h}^{d-3}}{4G_{N}}~{},$ (7)
$T=\frac{(d-1)r_{h}}{4\pi l^{2}}+\frac{(d-3)(1-\gamma^{2}l^{2}\mu^{2})}{4\pi r_{h}}~{}.$ (8)
The traditional way of deriving the Smarr formula is by applying Euler’s theorem for homogeneous functions with respect to the "bulk" scaling of thermodynamic quantities [^1] Scaling arguments have been extensively employed in the black hole thermodynamics literature to derive Smarr formulas, accommodating a range of matter contents and asymptotic behaviors (see, e.g., Lu ; Alfaro1 ). A more formal approach to obtaining the Smarr formula involves evaluating the Komar integral at spatial infinity and the horizon. For recent developments in this approach, see Tomas1 ; Tomas2 ; Tomas3 .. For the purposes of this article, it is better to derive the Smarr formula in a different way. Let’s consider the grand canonical potential $\Omega$:
$\Omega(T,\mu,V)=U-TS-\mu Q~{}.$ (9)
In our context, $M$ represents the internal energy $U$ of the system. It is important to note that, in principle, $\Omega$ is a function of the volume $V$, which, in the case of black holes in AdS spacetimes, corresponds to the spatial volume of the boundary metric where the dual field theory resides. In this work, we interpret black hole thermodynamics as the thermodynamics of the dual field theory in the strongly interacting regime, in line with the AdS/CFT correspondence. For the AdSd Reissner-Nordström black hole, we compute the grand canonical potential $\Omega$ using the Euclidean on-shell action
$\Omega=\frac{A_{d-2}}{8\pi G_{N}}r^{d-3}_{h}-\frac{M}{(d-2)}~{}.$ (10)
Equating the definition of $\Omega$ in (9) with the "on-shell value" of $\Omega$ (10), we obtain the Smarr formula for AdS black holes Karch :
$(d-1)M=(d-2)(TS+\mu Q)+\frac{(d-2)A_{d-2}}{8\pi G_{N}}r^{d-3}_{h}~{}.$ (11)
In this expression, the event horizon radius $r_{h}$ should be understood as a function of the thermodynamic variables. In summary, the Smarr formula is an unusual relation among thermodynamic quantities.
In this work, we aim to develop a deeper physical understanding of the Smarr formula within the broader framework of the AdS/CFT correspondence. We will demonstrate that the Smarr formula is a natural generalization of the Euler equation, $U+pV=TS+\mu Q$, with an additional correction due to a geometrical scale inherent to the physical problem. While a similar idea is explored in Karch , our approach extends beyond pure thermodynamic arguments by leveraging insights from fluid/gravity duality.
In thermodynamics, the Euler equation arises in the thermodynamic limit, where the volume $V$ and the number of particles $N$ both tend to infinity while maintaining a fixed ratio $\frac{N}{V}$. In this limit, the grand canonical potential becomes $\Omega=-pV$, and the Euler equation follows as $-pV=U-TS-\mu Q$. We seek to extend this reasoning to the Smarr formula. However, in the context of black hole thermodynamics, defining the thermodynamic volume $V$ and pressure $p$ is far from straightforward. This is where fluid/gravity duality offers crucial insight. To define the pressure $p$ for AdS black holes, we observe that the boundary stress-energy tensor associated with these black holes takes the form of a perfect fluid stress tensor Clifford0 ; Clifford . The boundary stress tensor for any given gravitational solution can be calculated using the formula provided in Kraus :
$T_{ab}=\frac{1}{8\pi G_{N}}\left(h_{ab}\left(L-\frac{(d-3)}{l}\right)+K_{ab}+\frac{lG_{ab}}{(d-4)}\right)~{},$ (12)
where $h_{ab}$ is the induced metric, $K_{ab}$ is the extrinsic curvature, and $G_{ab}$ is the Einstein tensor for a fixed value of the radial direction $r$. As a reminder, $d$ is the bulk dimension, so $(d-1)$ is the boundary dimension. In our case, the boundary metric is
$ds^{2}|_{\partial\mathcal{M}}=-dt^{2}+l^{2}d\Omega_{d-2}^{2}~{}.$ (13)
Thus, the volume of the spatial section of the boundary spacetime is $V=l^{d-2}A_{d-2}$ where the AdS radius $l$ is acting as the radius of the $(d-2)$-dimensional sphere. The components of the boundary stress tensor are
$T^{t}_{\;t}=\frac{(d-2)\tilde{m}}{16\pi G_{n}l^{2}}=\frac{M}{l^{d-2}A_{d-2}}=\frac{M}{V}~{},$ (14)
$T^{\phi_{j}}_{\;\phi_{j}}=\frac{\tilde{m}}{16\pi G_{n}l^{2}}=\frac{M}{(d-2)l^{d-2}A_{d-2}}=\frac{M}{(d-2)V}~{}.$ (15)
where we denoted by $\phi_{j}$ an arbitrary angle of the sphere. We can recast these components of $T_{ab}$ as a perfect fluid stress tensor
$T_{ab}=pg_{ab}+(p+\epsilon)u_{a}u_{b}~{},$ (16)
where $g_{ab}$ denotes the boundary metric, and $u_{a}$ is a unit timelike vector. The pressure $p$ and the energy density $\epsilon$ are
$\epsilon=\frac{M}{V}~{},\quad p=\frac{\epsilon}{(d-2)}~{}.$ (17)
This holographic perfect fluid in certain sense is dual to the AdSd Reissner-Nordström black hole. Besides, this fluid satisfies the conformal equation of state $\epsilon=(d-2)p$. The thermodynamics of this fluid is the same than the black hole thermodynamics since the fluid is at global thermal equilibrium.
Returning to the Smarr formula, substituting the conformal equation of state $M=(d-2)pV$ into (11) provides a more suggesting expression:
$M+pV=TS+\mu Q+\frac{A_{d-2}}{8\pi G_{N}}r^{d-3}_{h}~{},$ (18)
which resembles the Euler equation plus an additional correction term.
As mentioned earlier, the traditional Euler equation arises in the thermodynamic limit, where $V\to\infty$. However, in our case, the volume $V$ occupied by the fluid is finite, and thus the standard derivation of the Euler equation no longer applies. In the analysis leading to equation (18), we relied on the idea that a perfect fluid in global thermal equilibrium provides an equivalent description of global thermodynamics. Yet, perfect fluids are generally used to describe local thermodynamics, where the Euler equation is assumed to always hold—a point that appears contradictory to our intuitive understanding of the Smarr formula that we want to verify.
To fully comprehend the physical significance of the Smarr formula, it is necessary to generalize the Euler equation to apply in a broader setting. For this reason, in the next section, we will set aside the discussion of black holes and focus exclusively on the Euler equation in the context of perfect fluids.
3 Effective action for perfect fluids
In this section, we begin by reviewing the standard construction of the effective action for perfect fluids, which naturally leads to the Euler equation. We then extend this framework by introducing a new variable into the effective action, capturing the fluid’s response to the presence of a geometrical scale. This scale is relevant in certain spacetimes where the fluid resides. With this extension, we derive a generalized form of the Euler equation.
3.1 Review standard effective action for perfect fluids
Currently, there are formal constructions of effective actions for hydrodynamics using techniques such as the Schwinger-Keldysh formalism Liu2 ; Liu3 ; Ranga1 ; Ranga2 and superspace methods Ranga3 ; Ranga4 ; Jensen1 , with comprehensive reviews available in Liu4 ; Basar . Although these advanced techniques will not be explicitly employed in this article, they highlight the significant progress and depth this field has achieved. Instead, we revisit the foundational constructions presented in Dubovsky1 ; Son ; Dubovsky2 , informed by a more modern understanding of the subject. Moreover, we focus exclusively on perfect fluids, setting aside considerations of dissipation and hydrodynamic anomalies.
To construct the effective action for perfect fluids, two components are required:
  • Auxiliary Scalar Fields: To ensure that the covariant conservation of the stress tensor becomes a second-order equation (as it is initially a first-order equation with respect to the fluid variables), auxiliary scalar fields are introduced as the low-energy dynamical variables of the effective action.
  • Symmetry Organization: To organize the potential terms that can appear in the action principle, the symmetries of the fluid must be known. In this case, the relevant symmetries are low-energy gauge symmetries, which imply that the auxiliary scalar fields can be interpreted as Stueckelberg fields Liu2 ; Liu4 .
Given the dynamical variables and symmetries of the effective action, we should, in principle, consider all gauge-invariant terms in the theory. However, we will disregard higher-derivative terms in the effective action since our interest is limited to perfect fluids.
In the introduction, we assumed that the bulk spacetime has $d$ dimensions. Consequently, in this section, we will consider a perfect fluid residing in a $(d-1)$-dimensional curved spacetime, though our focus will be on the field-theoretical aspects of the discussion. The connection between the results of this section and gravity will be explored in the following section. Let’s begin by establishing the notation: the background metric for the perfect fluid is
$ds^{2}=g_{ab}dx^{a}dx^{b}~{},\quad a,b\in\{0,1,...,d-2\}~{}.$ (19)
The low-energy dynamical variables for the effective action are $(d-2)$ spatial scalar fields $\sigma^{I}$ that describe the fluid element, plus a temporal scalar field $\sigma^{0}$ that acts as an internal clock for the fluid element:
$\sigma^{I}=\sigma^{I}(x^{a})~{},\quad\sigma^{0}=\sigma^{0}(x^{a})~{},\quad I\in\{1,...,d-2\}~{}.$ (20)
The fluid dynamics possess the following emergent low-energy internal symmetries:
$\sigma^{0}\to\sigma^{0}+f^{0}(\sigma^{I})~{},$ (21)
$\sigma^{I}\to\sigma^{I}+f^{I}(\sigma^{J})~{}.$ (22)
These symmetries transformation can be motivated by noticing that $\sigma^{I}$ represents a spatial coordinates describing the fluids. Thus, $\sigma^{I}$ must satisfy the comoving condition
$\frac{d}{d\tau}\sigma^{I}(x^{a}(\tau))=0\Longleftrightarrow u^{a}\partial_{a}\sigma^{I}=0~{},$ (23)
where $x^{a}$ in the position in the spacetime, and the fluid velocity is $u^{a}=\frac{dx^{a}}{d\tau}$. In the second expression of the comoving condition, we use the directional derivative along the fluid velocity $u^{a}$. The comoving condition (23) is preserved by the symmetry transformations (21), (22). It is also important to note that the analogous expression for the temporal scalar field, $\frac{d}{d\tau}\sigma^{0}(x^{a}(\tau))$ is non-zero, but remains gauge-invariant under these symmetry transformations (21), (22).
In essence, the perfect fluid is described by a target metric $G_{MN}(\sigma^{M})$ with fluid coordinates $\sigma^{M}=(\sigma^{0},\sigma^{I})$. In this way, the internal symmetries (21), (22) are the isometries of degenerate metric Son
$ds^{2}=G_{MN}d\sigma^{M}d\sigma^{N}~{},\quad G_{NM}n^{M}=0~{},\quad M,N\in\{0,1,...,d-1\}~{},$ (24)
where $n^{N}$ is a null-like eigenvector associated with the zero eigenvalue of $G_{MN}$. An explicit parameterization of this degenerate target metric in terms of spatial target metric [^2] In the context of fluid/gravity duality, the target metric $G_{MN}$ is essentially the horizon metric. In this way, the horizon asymptotic symmetry becomes the emergent low-energy gauge symmetries of the effective action for hydrodynamics Liu5 . is
$ds^{2}=0\cdot(d\sigma^{0})^{2}+G_{IJ}(d\sigma^{I}-v^{I}d\sigma^{0})(d\sigma^{J}-v^{J}d\sigma^{0})~{},\quad I,J\in\{1,...,d-1\}~{},$ (25)
where $v^{J}$ denotes boosts. The null normal vector $n^{M}$ is then given by
$n^{M}=\frac{1}{\gamma}(1,v^{J})~{},\quad\gamma^{-1}=\sqrt{1-G_{IJ}v^{I}v^{J}}~{}.$ (26)
Thus, the fluid coordinates $\sigma^{M}$ are Stueckelberg fields associated with a sub-class of diffeomorphisms of the target metric. Further discussion of this point is found in Liu2 ; Liu4 .
So far, we have provided the description for neutral fluids. To describe charged fluids, a Stueckelberg field associated with a $U(1)$ symmetry must be introduced Dubovsky2
$\varphi=\varphi(x^{a})~{},$ (27)
with the following internal transformation, dubbed chemical shift symmetry:
$\varphi\to\varphi+g(\sigma^{I})~{}.$ (28)
Given the low-energy dynamical variables $\sigma^{a}$, $\varphi$, and the associated low-energy gauge symmetries, we can now construct the effective action for a charged perfect fluid. Using the auxiliary fluid variables $\sigma^{a}$ and $\varphi$, along with the fluid velocity $u^{a}$ and the target metric $G_{MN}$, we form diffeomorphism-invariant scalars with respect to the physical metric $g_{ab}$. These scalars must also be gauge-invariant under the fluid symmetry transformations (21) and (22). The first of these scalars, the entropy density $s$ of the system, is defined in terms of the spatial target metric $G_{IJ}$ (25) as follows Son :
$s=\sqrt{\text{Det}(B^{IJ}G_{JK})}~{},\quad B^{IJ}=g^{ab}e_{a}^{I}e_{b}^{J}~{},$ (29)
$e_{a}^{I}\equiv\partial_{a}\sigma^{I}-v^{I}\partial_{a}\sigma^{0}~{}.$ (30)
Another scalar is the chemical potential $\mu$ associated with the $U(1)$ symmetries Dubovsky2 which reads
$\mu=u^{a}\partial_{a}\varphi~{}.$ (31)
We note that both $s$ and $\mu$ are local functions of the physical spacetime where fluids is on, and they take constant values only at global thermal equilibrium.
The low-energy fields $\sigma^{M}$, $\varphi$ are auxiliary variables that allow the description of perfect fluids and their local thermodynamics in terms of an action principle. To zero-order in the derivative expansion, the effective action for a perfect fluid is:
$I=-\int d^{d-1}x\sqrt{-g}\zeta(s,\mu)~{},$ (32)
where $\zeta$ is a local thermodynamic potential function of the entropy density $s$ and the chemical potential $\mu$. By taking the variation with respect to the background metric $g_{ab}$, one obtains the stress tensor:
$T_{ab}=\left(-\zeta+s\frac{\partial\zeta}{\partial s}\right)g_{ab}+\left(-\mu\frac{\partial\zeta}{\partial\mu}+s\frac{\partial\zeta}{\partial s}\right)u_{a}u_{b}~{}.$ (33)
Comparing this result with the stress tensor of a perfect fluid $T_{ab}=pg_{ab}+(p+\epsilon)u_{a}u_{b}$, one deduces the pressure $p$
$p=-\zeta+s\frac{\partial\zeta}{\partial s}~{},$ (34)
and the local energy density $\epsilon$
$\epsilon=\zeta-\mu\frac{\partial\zeta}{\partial\mu}~{}.$ (35)
The local thermodynamic potential $\zeta$ is essentially a Legendre transformation: $\zeta=\epsilon-\mu q$, where the charge density is $q=-\frac{\partial\zeta}{\partial\mu}$. Then, the temperature is simply $T=-\frac{\partial\zeta}{\partial s}$. In this way, the consistency of the effective action (32) and the stress tensor (33) with the standard thermodynamic relation implies the Euler equation:
$p+\epsilon=sT+\mu q~{}.$ (36)
Thus, the Euler equation is an unavoidable consequence of the effective action (32). Note that the equation of state for the fluid is not used in this derivation.
For completeness, we mention that due to the comoving condition, the entropy current $s^{a}$ is conserved off-shell:
$\nabla_{a}s^{a}=0,\quad s^{a}=su^{a}~{}.$ (37)
This conservation law reflects the non-dissipative nature of the fluid. In addition, the chemical shift symmetry (28) implies the on-shell conservation of the charge current:
$\nabla_{a}q^{a}=0,\quad q^{a}=qu^{a}~{}.$ (38)
With this, we conclude our review of the effective action formulation for perfect fluids. In the following section, we will use this framework to derive a generalized Euler equation.
3.2 Generalized Euler equation
To generalize the Euler equation, the only viable approach is to introduce a new variable into the effective action for perfect fluids. This new variable must be a diffeomorphism-invariant scalar with respect to $g_{ab}$ and gauge-invariant under the symmetry transformations defined in (21) and (22). As discussed at the end of section 2, there are situations where the standard scaling argument for deriving the Euler equation breaks down—such as when the system’s volume is compact. In this case, the radius of the compact space introduces an additional geometrical scale to the physical problem. This suggests that the new variable in the effective action should have a geometrical origin. The gauge-invariant quantity that satisfies these requirements is
$y=u^{a}\partial_{a}\sigma^{0}~{}.$ (39)
The definition of this variable closely resembles the definition of the chemical potential $\mu$ presented in (31). However, the temporal scalar field $\sigma^{0}$, which represents the clock for the fluid element, differs in nature from the Stueckelberg field $\varphi$ associated with U(1) symmetry [^3] The variable $y$ has been previously studied in the literature, but Ballesteros did not clearly distinguish between the temporal scalar field $\sigma^{0}$ and the gauge scalar field $\varphi$.. Given the definition of $y$, its significance may not be immediately clear. Therefore, let us first derive the generalized Euler equation and subsequently discuss the physical meaning of $y$.
The effective action for perfect fluids is now
$I=-\int d^{d-1}x\sqrt{-g}\chi(s,\mu,y)~{},$ (40)
where $\chi$ represents a new local thermodynamic potential that depends on entropy density $s$, chemical potential $\mu$, and the new variable $y$. Taking the variation with respect to the background metric $g_{ab}$, the stress tensor follows
$T_{ab}=\left(-\chi+s\frac{\partial\chi}{\partial s}\right)g_{ab}+\left(-y\frac{\partial\chi}{\partial y}-\mu\frac{\partial\chi}{\partial\mu}+s\frac{\partial\chi}{\partial s}\right)u_{a}u_{b}~{}.$ (41)
The next step is to compare this expression with the stress tensor of a perfect fluid: $T_{ab}=pg_{ab}+(p+\epsilon)u_{a}u_{b}$. Thus, we can identify the pressure $p$ and energy density $\epsilon$ as:
$p=-\chi+s\frac{\partial\chi}{\partial s}~{},\quad\epsilon=\chi-y\frac{\partial\chi}{\partial y}-\mu\frac{\partial\chi}{\partial\mu}~{}.$ (42)
The following combination will be relevant later:
$\epsilon+p=-y\frac{\partial\chi}{\partial y}-\mu\frac{\partial\chi}{\partial\mu}+s\frac{\partial\chi}{\partial s}~{}.$ (43)
As a second step, we need to properly identify the rest of the thermodynamic variables. We note that the energy density has the following dependence: $\epsilon=\epsilon(s,q,y)$. Thus, if we perform the Legendre transformation $\tilde{\epsilon}=\epsilon-\mu q$, we find that $\tilde{\epsilon}=\tilde{\epsilon}(s,\mu,y)$. In this way, the temperature $T$ and charge density $q$ are given by:
$T=\left.\frac{\partial\tilde{\epsilon}}{\partial s}\right|_{y,\mu}~{},\quad q=\left.\frac{\partial\tilde{\epsilon}}{\partial\mu}\right|_{y,s}~{}.$ (44)
From these expressions, we deduce how to compute the local temperature $T$ and charge density $q$ from the thermodynamic potential $\chi$:
$T=\frac{\partial\chi}{\partial s}-y\frac{\partial^{2}\chi}{\partial y\partial s}~{},\quad q=-\frac{\partial\chi}{\partial\mu}~{}.$ (45)
Inserting these results into (43), we obtain the relation
$\epsilon+p=sT+\mu q+y\partial_{y}\left(-\chi+s\frac{\partial\chi}{\partial y}\right)~{}.$ (46)
Inside the parenthesis of the last term, we recognize the local pressure $p$ given in (42). Therefore, we obtain a new relation among thermodynamic variables:
$\epsilon+p=sT+\mu q+y\frac{\partial p}{\partial y}~{}.$ (47)
This expression is the generalized Euler equation that we were seeking. The additional correction involves a derivative of the pressure $p$ with respect to $y$. Therefore, if the thermodynamics of a fluid in a given spacetime does not respond to variations in $y$, the extra correction $\frac{\partial p}{\partial y}$ vanishes. Another equivalent way of writing (47), which may be useful for clarity, is as follows:
$\omega=\epsilon-sT-\mu q\Longrightarrow\omega=-p+y\frac{\partial p}{\partial y}~{},$ (48)
where $\omega$ is the density of the grand canonical potential. This form clearly illustrates that the additional scale in the problem, parameterized by $y$, contradicts the conventional wisdom that, in the thermodynamic limit, the grand canonical potential $\omega$ is proportional to the pressure $p$.
To understand the meaning of $y$, let us consider a scenario of global thermal equilibrium, where it is possible to align the fluid clock with physical time, i.e., $\sigma^{0}=x^{0}$. In this case, we obtain $y=u^{0}$, an expression that may be challenging to interpret. To gain some intuition, let us perform a Weyl transformation of the physical metric $g^{\prime}=Wg$, with $W$ being a conformal factor. Under this transformation, the fluid velocity changes as $u^{\prime a}=\frac{u^{a}}{W}$. Consequently, we arrive at the following relation:
$y=u^{\prime 0}=\frac{u^{0}}{W}~{}.$ (49)
This relation suggests that the variable $y$ is associated with the inverse of the conformal factor $W$. Instead of performing an explicit Weyl transformation, we can express the physical metric $g$ as a conformal factor times a reference metric $\bar{g}$, i.e., $g=W\bar{g}$.
It is important to emphasize that the motivation for introducing the variable $y$ was to capture an additional scale that a given spacetime could introduce into the thermodynamics of a fluid. This is why, we should not expect a microscopic interpretation for the variable $y$, unlike the other variables in the new thermodynamic potential $\chi=\chi(s,\mu,y)$.
4 Smarr formula as a particular case of the generalized Euler equation
As discussed around equation (18), the Smarr formula can be viewed as a generalization of the Euler equation. Traditionally, however, the thermodynamics of black holes is examined using Euclidean methods. Unfortunately, these methods do not readily define the thermodynamic volume $V$ and pressure $p$ of a thermal system within the context of gravity. This limitation motivates our use of fluid dynamics to analyze the Smarr formula. We will now demonstrate the generalized Euler equation (47), derived in the previous section, reduces to a local version of the Smarr formula when applied to holographic fluids. Subsequently, the standard form of the Smarr formula emerges by considering the fluid in global thermal equilibrium, which encapsulates the same thermal information as the first law of thermodynamics.
In this section, we extend the discussion from section 2 by examining the AdSd Reissner-Nordström black hole with arbitrary horizon topology. We adopt Eddington-Finkelstein coordinates, as this gauge simplifies the analysis of Weyl transformations compared to the Fefferman-Graham gauge. Moreover, we restrict the boundary dimension $(d-1)$ to be odd; otherwise, the conformal anomaly would need to be accounted for in the effective action. The line element is:
$ds^{2}=-f(r)du^{2}-2dudr+r^{2}d\Sigma_{k}^{2}~{},\quad f(r)=\frac{r^{2}}{l^{2}}+k-\frac{\tilde{m}}{r^{d-3}}+\frac{\tilde{q}^{2}}{r^{2d-6}}~{}.$ (50)
Here, the gauge field $A$ is still given by equation (4). The constant $k$ represents the unit-normalized curvature of the transverse space $\Sigma_{k}$. Thus, the black brane is recovered for $k=0$, the spherical black hole for $k=1$, and the hyperbolic black hole for $k=-1$. In all cases, the black hole energy $M$ and charge $Q$ are related to the parameters $\tilde{m}$ and $\tilde{q}$ as follows:
$M=\frac{(d-2)\Sigma_{k,d-2}\tilde{m}}{16\pi G_{N}}~{},\quad Q=\frac{\gamma l(d-2)\Sigma_{k,d-2}\tilde{q}}{8\pi G_{N}}~{}.$ (51)
The entropy of the black holes is given by:
$S=\frac{\Sigma_{k,d-2}r_{h}^{d-3}}{4G_{N}}~{},$ (52)
where $\Sigma_{k,d-2}$ represents the volume of the transverse space with unit radius. It is important to note that for flat and hyperbolic spaces, $\Sigma_{k,d-2}$ is formally infinite. Additionally, hyperbolic space has negative curvature, introducing an extra scale into the physical problem.
The Hawking temperature $T$ and the chemical potential $\mu$ are
$T=\frac{(d-1)r_{h}}{4\pi l^{2}}+\frac{(d-3)(k-l^{2}\gamma^{2}\mu^{2})}{4\pi r_{h}}~{},\quad\mu=\frac{\tilde{q}}{\gamma lr_{h}^{d-3}}~{}.$ (53)
Note that the chemical potential $\mu$ is the same as in equation (5). The Smarr formula is now given by:
$M+pV=TS+\mu Q+\frac{kA_{d-2}}{8\pi G_{N}}r^{d-3}_{h}~{}.$ (54)
For for $k=0$ the last term vanishes, thereby recovering the standard Euler equation. This is because, in flat space, there are no additional scales to consider.
We will now compute the thermodynamic potential $\chi$ (40) which is associated with the effective action for the holographic perfect fluid dual to the AdSd Reissner-Nordström black hole. Our objective is to place the holographic fluid in a boundary spacetime that exhibits spacetime dependence, allowing us to mimic a state that is near out of equilibrium. As a consequence, the thermodynamic variables now will become local local functions of the spacetime. To obtain the boundary metric with the desired properties, we perform the following radial diffeomorphism in the bulk line element (50):
$r\to W(x^{a})r~{},$ (55)
where $W(x^{a})$ is an arbitrary function independent of the radial variable $r$. The new line element is
$ds^{2}=-f(Wr)du^{2}-2Wdudr-2r\partial_{a}Wdudx^{a}+W^{2}r^{2}d\Sigma_{k}^{2}~{},$ (56)
$f(r)=\frac{W^{2}r^{2}}{l^{2}}+k-\frac{\tilde{m}}{(Wr)^{d-3}}+\frac{\tilde{q}^{2}}{(Wr)^{2d-6}}~{}.$ (57)
Therefore, the boundary metric is
$ds^{2}|_{\partial\mathcal{M}}=-\frac{W^{2}}{l^{2}}du^{2}+W^{2}d\Sigma_{k}^{2}~{}.$ (58)
This line element describes a non-homogeneous, time-dependent background in which the fluid resides. The radial diffeomorphism (55) consequently induces a Weyl transformation on the boundary metric, leading to the following expression for the fluid velocity:
$u_{a}=\frac{W}{l}\delta^{a}_{0}~{}.$ (59)
Let’s focus now on the target metric $G_{MN}$ that describes the fluid space. Using the unitary gauge embedding $\sigma^{0}=x^{0}$ and $\sigma^{I}=x^{i}$ Son , the target metric $G_{MN}$ takes a very simple form:
$ds^{2}_{fluid}=-0\cdot du^{2}+r^{2}_{+}d\Sigma^{2}_{k}~{},$ (60)
that matches the horizon metric (with no conformal factor). Given the target metric and the fluid velocity, we can use the formulas (29), (31), and (39) to compute the entropy density $s$, the chemical potential $\mu$, and the variable $y$, respectively:
$s=\frac{\Sigma_{k,d-2}}{4G_{N}}\frac{r_{h}^{d-2}}{W^{d-2}}~{},\quad\mu=\frac{\tilde{q}}{\gamma lr_{h}^{d-3}}\frac{1}{W}~{},\quad y=\frac{l}{W}~{},$ (61)
which are the variables that the thermodynamic potential $\chi$ depends on. We will utilize these variables to derive the explicit expression for the new thermodynamic potential $\chi$ from the energy density $\epsilon$ using equation (42). Our next step involves computing the boundary stress tensor via the formula (12) and the boundary metric given in (58). The components of the stress tensor are
$T^{t}_{\;t}=\frac{M}{W^{d-1}\Sigma_{k,d-2}}~{},$ (62)
$T^{\phi_{j}}_{\;\phi_{j}}=\frac{M}{(d-2)W^{d-1}\Sigma_{k,d-2}}~{}.$ (63)
Similar to section 2, these components can be expressed in the form of a perfect fluid stress tensor $T_{ab}=pg_{ab}+(p+\epsilon)u_{a}u_{b}$, from which we read the energy density $\epsilon$:
$\epsilon=\frac{(d-2)\Sigma_{k,d-2}}{16\pi G_{N}l^{2}}\left[\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l^{2})^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{},$ (64)
where $M$ has been expressed in terms of $r_{h}$ as an intermediate step, and then in terms of the variables $s$, $\mu$, and $y$ given in equation (61). One can verify that the conformal equation of state: $p=\frac{\epsilon}{d-2}$ is satisfied.
As a final step, using equation (42) that connects $\epsilon$ and $p$ with the thermodynamic potential $\chi$, we can infer the effective action for this holographic fluid
$I=-\int d^{d-1}x\sqrt{-g}\chi(s,\mu,y)~{},$ (65)
with the thermodynamic potential $\chi$ given by
$\chi=\frac{(d-2)\Sigma_{k,d-2}}{16\pi G_{N}l^{2}}\left[-\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l^{2})^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{}.$ (66)
We note that all thermodynamic variables involved in $\chi$ are local functions of the spacetime where the fluid resides. Certainly, we have derived $\chi$ in this case through inverse engineering, as we already knew some of the thermodynamic variables and the geometrical variable $y$. This approach is not an inconvenience for us; rather, it allows us to treat the thermodynamic potential $\chi$ as a starting point, as it encodes all the thermal information of the fluid. Finally, using equation (45), we compute respectively the local temperature $T$ and charge density $q$ as follows:
$T=\frac{(d-1)}{4\pi l^{2}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{1}{d-2}}+\frac{(d-3)(ky^{2}-(\gamma l^{2})^{2}\mu^{2})}{4\pi l^{2}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{-1}{d-2}}~{},$ (67)
$q=\frac{(d-2)\Sigma_{k,d-2}\gamma^{2}l^{2}\mu}{8\pi G_{N}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (68)
Furthermore, we notice that $\frac{\partial p}{\partial y}$ is proportional to $k$:
$\frac{\partial p}{\partial y}=k\frac{\Sigma_{k,d-2}y}{8\pi G_{N}l^{2}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (69)
Thus, this term vanishes for a charged black brane, which is an expected result due to the absence of any extra geometrical scale arising from flat spacetime. By performing some straightforward algebra, we can verify that the generalized Euler equation holds:
$\epsilon+p=sT+\mu q+y\frac{\partial p}{\partial y}~{}.$ (70)
To this point, we have utilized an arbitrary conformal factor to represent a curved physical background, simulating a state near out of equilibrium for the fluid. The usual global thermal state is obtained by selecting $W=Rl$, where $R$ is a constant factor, reducing equation (70) to the Smarr formula given in (54). Strictly speaking, the last statement is true for $R=1$; however, the extra scale allows us to distinguish the radius $Rl$ of the transverse space $\Sigma_{k}$ from the AdS radius $l$.
In conclusion, we have demonstrated that the Smarr formula is a particular instance of the generalized Euler equation derived in section 3.2. This finding suggests that holographic fluids adhering to the Smarr formula are not fundamentally different from other more conventional fluids. The key takeaway is that when an extra scale emerges from the geometry in which the fluid or thermal system resides, the standard scaling argument used to derive the Euler equation no longer applies, necessitating a generalization of this equation. This generalization was provided in section 3.2, where we derived the generalized Euler equation (70). In cases with a compact volume, the correction term $y\frac{\partial p}{\partial y}$ in the Euler equation can be interpreted as a finite-size effect.
5 Comments on extended black hole thermodynamics
In the previous sections, we argued and demonstrated that the Smarr formula is a specific case of the generalized Euler equation, particularly in the context of black holes in AdS. In this sense, the Smarr formula naturally fits within the framework of the AdS/CFT correspondence, with its origins rooted in well-established physics. This conclusion is supported by the fact that we derived the generalized Euler equation using effective field theory techniques applied to fluids. Only after establishing this result did we invoke the AdS/CFT correspondence to interpret the Smarr formula.
Extended black hole thermodynamics was initially proposed outside the context of gauge/gravity duality to provide an interpretation of the Smarr formula. In the original version of this proposal, the cosmological constant $\Lambda$ was identified with a thermodynamic pressure Kastor ; Gibbons ; Dolan1 ; Dolan2 ; Dolan3 as follows:
$p_{\Lambda}=-\frac{\Lambda}{8\pi G_{N}}=\frac{(d-1)(d-2)}{16\pi G_{N}l^{2}}~{},$ (71)
where $d$ is the dimension of the bulk spacetime [^4] For reviews about extended black hole thermodynamics see Mann4 ; Mann5 . Other relevant literature in the subject Mann2 ; Mann3 ; Clifford2 .. Our understanding of the results in Karch suggests that this identification is problematic within the context of holography.
Following the work of Karch , a revised version of extended black hole thermodynamics was proposed to reinterpret the Smarr formula consistently with the AdS/CFT correspondence. To understand this new proposal, we recall the holographic identification between the parameters of gravity and those of the dual field theory:
$N^{2}\sim\frac{l^{d-2}}{G_{N}}~{}.$ (72)
Here, $N$ represents the number of degrees of freedom on field theory side of the duality. For instance, in $\mathcal{N}=4$ SYM, $N$ corresponds to the rank of the gauge group $SU(N)$, which is also the size of a matrix $A$ in the adjoint representation. In the revised version of extended black hole thermodynamics, the so-called central charge $C\sim N^{2}$ is treated as a thermodynamic variable associated with the cosmological constant $\Lambda$ Visser1 ; Gao ; Visser2 ; Mann6 . In this view, a variation in the cosmological constant induces a flow in theory space, affecting the thermodynamics of black holes in AdS. This proposal is supported by a possible origin of extended black hole thermodynamics in higher-dimensional physics Pedraza .
The main result of this paper—that the Smarr formula is a specific instance of the generalized Euler equation—contradicts this new proposal for extended black hole thermodynamics. The reasoning is as follows: perfect fluids, a particular case of hydrodynamics, differs fundamentally from renormalization group (RG) physics. While the number of degrees of freedom $N$ can indeed change with the energy scale in RG flow (as is well understood in the context of AdS/CFT, see Ioannis1 ; Ioannis2 ), this cannot explain the Smarr formula. In a perfect fluids, thermodynamic variables must be functions of spacetime to describe local thermal equilibrium, something that the number of degrees of freedom $N$ cannot do. The point is that we should not confuse a flow in a physical space where the fluid is on with a flow in theory space where an RG flow takes place. Finally, it is crucial to recognize that global equilibrium is merely a special case of local equilibrium. As a result, there is no justification for treating $C\sim N^{2}$ as a thermodynamic variable.
As a final remark for this section, the standard techniques for studying AdS black hole thermodynamics—such as the Euclidean method or your preferred method for computing conserved charges—struggle to clarify the meaning of the Smarr formula. One reason for this difficulty is the double role played by the AdS radius $l$. On one hand, $l$ can be interpreted as the radius of the transverse space at the boundary, while on the other, it sets the scale for constructing dimensionless quantities in holography, such as in equation (72). This point can be a source of confusion. However, in our construction, this issue is resolved because the variable $y$ (as defined in equation (39) is diffeomorphism-invariant with respect to physical metric $g_{ab}$ and gauge-invariant under the fluid transformations (21), (22) from the outset. Thus, there is no ambiguity. Ultimately, the insights gained from fluid/gravity duality were instrumental in revealing the correct understanding of the Smarr formula within holography.
6 Conclusions and future directions
In this work, we derived a generalized Euler equation (47) using effective field theory techniques, granting our result a broader range of applicability independent of the AdS/CFT correspondence. We then applied the generalized Euler equation to a specific case: a holographic fluid derived from the boundary stress tensor associated with an AdSd Reissner-Nordström black hole, from which we recovered the Smarr formula. Based on these findings, we further argued that the framework of extended black hole thermodynamics cannot be physically correct. This is because any new thermodynamic variable that one might attempt to associate with the AdS radius $l$ cannot become a function of the boundary spacetime, as required for local equilibrium.
It has been demonstrated that evaluating the on-shell action of spacetimes with a future event horizon (featuring a planar horizon topology) can reproduce the standard effective action presented in (32). This was achieved by setting up a double-Dirichlet boundary condition to evaluate the on-shell action Natalia1 ; Liu1 . It would be intriguing to extend this construction to the more general effective action (40) discussed in our article. The primary challenge lies in the fact that the new thermodynamic potential $\chi=\chi(s,\mu,y)$ does not coincide with the grand canonical potential $\omega=\omega(T,\mu,y)$, as might typically be expected from the evaluation of the Euclidean on-shell action.
Another intriguing avenue for exploration would be the explicit construction of the thermodynamic potential $\chi$ for rotating black holes, which also exhibit a boundary stress tensor with a perfect fluid form Clifford0 ; Clifford . In Section 4, we focused on odd bulk dimensions $d$ to sidestep the complications of conformal anomalies that arise when the fluid resides in curved spacetimes with even boundary dimensions $(d-1)$. It would be valuable to develop the most general effective action for perfect fluids that can incorporate conformal anomalies. An example of this is the Schwarzian action, which can be interpreted as an effective action for an anomalous perfect fluid in 1D Jensen2 . A similar result applies to the boundary action of a non-rotating BTZ black hole with boundary sources turned on Jensen3 ; Kevin .
The Smarr formula remains valid in gravity theories with higher-order curvature corrections Tomas1 . This raises an important question: can the generalized Euler equation (47) accommodate these theories? If not, it would suggest that the Smarr formula once again points to the need for further generalizations of the Euler equation. In such a scenario, it is unclear to the author what the nature of the new variables introduced in the effective action might be. This uncertainty arises because the variable $y$ in (39) had a clear geometric motivation, making it challenging to envision new variables with equally well-defined physical interpretations.
The isometries of the target metric $G_{MN}$, as given in (24), correspond to the emergent low-energy gauge symmetries of the effective action for hydrodynamics. More broadly, horizon asymptotic symmetries ASG1 ; ASG2 ; ASG3 ; ASG4 ; ASG5 can be interpreted as these emergent low-energy gauge symmetries in the boundary theory. Therefore, any generalization of horizon symmetries implies the existence of new gauge symmetries for holographic effective actions. Investigating the connection between this work and the new symmetries identified in Liu5 would be an intriguing direction for future research.
Similarly, the isometries of the degenerate target metric $G_{MN}$ given in (24) are closely related to Carrollian diffeomorphisms. Motivated by this, it should be possible to formulate an effective action for Carrollian fluids CR1 ; CR2 ; CR3 ; CR4 ; CR5 ; CR6 ; CR7 ; CR8 . In this context, Carrollian diffeomorphisms would serve both as gauge symmetries of the effective action and as spacetime symmetries of the fluid’s background [^5] This perspective aligns naturally with the Gauge/Gravity duality Liu5 , as it likely follows from taking a flat limit of the fluid. There is also a body of work exploring the connection between Carrollian fluids and horizon symmetries CR9 ; CR10 ; CR11 ; CR12 ; CR13 .. With this action principle in place, the Smarr formula for asymptotically flat black holes should follow in a manner similar to the derivation presented in this article.
Acknowledgements.
R.M. would like to thank David Berenstein, Alejandra Castro, and Ioannis Papadimitriou for various discussions on related topics that, in part, inspired this project. R.M. would also like to thank Clifford Johnson and Jorge Zanelli for helpful discussions, as well as David Berenstein, Alejandra Castro, Adolfo Holguin, Sean McBride, Andrew Svesko, and Manus Visser for their comments on the manuscript. R.M. is grateful to the Centro de Estudios Científicos, Valdivia, and Pontificia Universidad Católica de Chile, Santiago, for their hospitality while part of this manuscript was being written. R.M.’s research was supported in part by the Department of Energy under grant DE-SC 0011702, and by the Chilean Fulbright Commission and ANID through the Beca Igualdad de Oportunidades ##\## 56180016.